OpenStudy (staceyg):

Someone please help!!! What is the slope of a line that passes through the pair of points (2,5) and (8,3).

OpenStudy (staceyg):

1/3, -1/3, 3, -3

OpenStudy (staceyg):

@tkhunny

OpenStudy (tkhunny):

You have to know how to calculate slope. It is fundamental and it WILL be on the exam. \(Slope = \dfrac{\Delta\;y}{\Delta\;x} = \dfrac{Change\;in\;y}{Change\;in\;x} = \dfrac{Rise}{Run} = \dfrac{y_{1}-y_{0}}{x_{1}-x_{0}}\) Something in there needs to speak to you.

OpenStudy (staceyg):

So rise/run. How do I find out the slope using that?

OpenStudy (staceyg):

I divide x by y ?

OpenStudy (staceyg):

@Embryo

OpenStudy (anonymous):

So the formula that tkhunny gave you is completely correct, but something that would help you out would be to answer a question you might have which is "what the heck are"\[y_1,y_0,x_1,x_0\]alrighty, to answer this, you take the points that you want to find the slope of and you label them, it doesn't matter which point you use as which, so you could do this, set\[(x_0,y_0)=(2,5)\]and then set your other point as,\[(x_1,y_1)=(8,3)\]then you just need to plug that into the stuff that tkhunny gave you, hopefully that clarifies things for you